OpenStudy (anonymous):

you have two dice. what is the probabilitly that you will roll a three on the first die and a three on the second?

OpenStudy (anonymous):

Well, the chance on the first dice is: \[\frac{ 1 }{ 6 }\] and same with the second dice: so \[\frac{ 1 }{ 6 }*\frac{ 1 }{ 6 } = \frac{ 1 }{ 6 }\] So you have a \[\frac{ 1 }{ 36 }\] chance to roll double threes.